I am not sure how or if there is a solution for this:
solve for y: 1/2y + 2 = 1/6(3y-9)

Answer by lildude_snowdog(5)

y=-.5
the 1/6*3y 1/6*9
1/2y+2=.5y-1.5
-.5y -.5y
y+2=1.5
-2 -2
y=-.5